OpenStudy (anonymous):

find two consecutive integers whose sum is 101

OpenStudy (anonymous):

50+51

OpenStudy (pottersheep):

So let the first number be x second is (x+1) x + (x + 1) = 101 2x = 100 x = 50 (the first number) so the second number is (x+1) so (50+1) = 51 Hope that helps :)